OpenStudy (anonymous):

Help!! will give medal :) the perimeter of a television screen is 150 inches. The ratio of the length to the width is 16 to 9. Find the length and width

OpenStudy (zzr0ck3r):

perimeter of tv 2lngth+2height we know 16/9 = length/height

OpenStudy (zzr0ck3r):

and we know perimeter = 150

OpenStudy (zzr0ck3r):

so 2l+2h=150 16/9 = l/h solve the second equation for l and tell me what you have

OpenStudy (zzr0ck3r):

it says width and not height?

OpenStudy (anonymous):

correct so i would do 2(16)+2(9)=150??

OpenStudy (zzr0ck3r):

no you need to solve for either l or h in the equation 16/9 = l/h so l = h(16/9) now plug that in for l in the first equation 2(16/9) + 2h = 150

OpenStudy (zzr0ck3r):

I should have used w not h, but its the same thing

OpenStudy (anonymous):

is that 16 divided by 9 or do u mean pick eather 16 or 9

OpenStudy (zzr0ck3r):

\[2\frac{16}{9}+2h=150\]

OpenStudy (anonymous):

how do i type that into my calculator??

OpenStudy (aaronq):

you have to isolate one of the variables and plug it into the other formula 2l+2h=150 ->l= 75-2h 16/9 = l/h -> 16/9=(75-2h)/h

OpenStudy (zzr0ck3r):

the way I did it the first time

OpenStudy (zzr0ck3r):

solve(2(16/9)+2x=150,x)

OpenStudy (zzr0ck3r):

depending on the calculator or 2(16/9) + 2h = 150 multiple everything by 9 32+18h = 1350 18h = 1350-32 18h = 1318 h=?

OpenStudy (zzr0ck3r):

find h, then pluf that back into either equation to find l

OpenStudy (zzr0ck3r):

pluf = plug

OpenStudy (anonymous):

ok i think i got it thanks for the help :) ill let you kno what i get

OpenStudy (anonymous):

i get 73.22 hight but would that be width @zzr0ck3r

OpenStudy (anonymous):

then would i do 2(16/9)2(73.22)=150

OpenStudy (zzr0ck3r):

no 2l+ 2(73.22) = 150 solve for l or 16/9 = l/73.22 solve for l

OpenStudy (anonymous):

i get 1.78 but that doesn't seem right

OpenStudy (anonymous):

i did 2l+146.44=150 2l=3.56 l=1.78

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

16/9 = l/73.22 l = (73.22)(16/9)

OpenStudy (zzr0ck3r):

I think you need to go back and work on algebra before you will be able to answer questions like this

OpenStudy (anonymous):

this is my last class and i haven't taken algebra in 5 er 6 years